OpenStudy (anonymous):

How can I solve this equation a factor. y = 3/5x - 3 2x - y = -4

OpenStudy (saifoo.khan):

a factor??

OpenStudy (anonymous):

y=2x+4 2x+4=3/5x -3 7/5x=-7 x=-5

OpenStudy (anonymous):

the answer is -5, -6. I just can't figure out how they got that.

OpenStudy (anonymous):

I just showed it.

OpenStudy (anonymous):

y=2x+4

OpenStudy (anonymous):

2x+4=3/5x -3

OpenStudy (saifoo.khan):

substituion!!!

OpenStudy (anonymous):

7/5x= -7

OpenStudy (anonymous):

x=-5

OpenStudy (anonymous):

Ok how did you get that. 2x+4=3/5x-3. How did you get 7/5= -7

OpenStudy (anonymous):

It's a secret.

OpenStudy (anonymous):

5y=3/5x-3 If I multiply both side by 5 I get 5y = 5(3/5)x - 5(3) 5y = 3x - 15

OpenStudy (anonymous):

x=-17/25 y=66/25